OpenStudy (anonymous):

Two numbers are said to be coprime if their highest common factor is 1. Prove that any two consecutive positive integers are always coprime.

OpenStudy (anonymous):

give me a min

OpenStudy (anonymous):

let n,p,q be integers , such that p>q>0 and n>1 then n*q and n*p are integers and are co-prime Let's assume that it's possible they are consecutive positive integers - so: n*q - n*p = 1 n(q-p)=1 q-p=1/n but, 1/n is not an integer , so q and p cannot both be integers- contradicting our assumption.

OpenStudy (anonymous):

n*q and n*p are not coprime... they have n as a common factor and n>1, so their highest common factor cannot be 1.

OpenStudy (anonymous):

Doesn't this sort of automatically follow from the division algorithm?

OpenStudy (anonymous):

OOOps.. I misread what co-prime means, and just proved something else then ...

OpenStudy (anonymous):

Please elaborate estudier.

OpenStudy (anonymous):

how about something along these lines :) Let p,n be integers and p be a factor of both n and (n+1) so: n/p is an integer (n+1)/p is an integer (n+1)/p=n/p+1/p = integer + 1/p integer + 1/p --> p=1

OpenStudy (anonymous):

p=1 is the only case where: integer + 1/p will be an integer as well.

OpenStudy (anonymous):

Good! A simpler proof is using the Euclidean algorithm to find the highest common factor of two numbers. gcd(n+1,n) = gcd(n,1) = 1 follows immediately like this! Thanks though!

OpenStudy (anonymous):

Thanks, that Euclidean algorithm rocks !